### v1

# MixNet: Towards Effective and Efficient UHD Low-Light Image Enhancement

Abstract
With the continuous advancement of imaging devices, the prevalence of Ultra-High-Definition (UHD) images is rising. Although many image restoration methods have achieved promising results, they are not directly applicable to UHD images on devices with limited computational resources due to the inherently high computational complexity of UHD images. In this paper, we focus on the task of low-light image enhancement (LLIE) and propose a novel LLIE method called MixNet, which is designed explicitly for UHD images. To capture the long-range dependency of features without introducing excessive computational complexity, we present the Global Feature Modulation Layer (GFML). GFML associates features from different views by permuting the feature maps, enabling efficient modeling of long-range dependency. In addition, we also design the Local Feature Modulation Layer (LFML) and Feed-forward Layer (FFL) to capture local features and transform features into a compact representation. This way, our MixNet achieves effective LLIE with few model parameters and low computational complexity. We conducted extensive experiments on both synthetic and real-world datasets, and the comprehensive results demonstrate that our proposed method surpasses the performance of current state-of-the-art methods. The code will be available at https://github.com/zzr-idam/MixNet.
1 Introduction
In recent years, high-resolution imaging has experienced significant advancements due to the emergence of sophisticated imaging sensors and displays. This has led to a rapid evolution of Ultra-High-Definition (UHD) imaging technology. However, the inherent high-resolution characteristics of UHD images make them more susceptible to noise during the imaging process. Additionally, the increased number of pixels in UHD images challenges the efficiency of existing image processing methods. These undoubtedly affect the quality of UHD imaging.
In this paper, we focus on the task of UHD low-light image enhancement (LLIE). Images captured under low-light conditions usually exhibit significant degradation, including reduced visibility, low contrast, and high noise levels. Recovering ideal normal-light images from UHD low-light images is a long-studied problem.
Figure 1: Model complexity and performance comparison between our proposed MixNet and other LLIE methods on the UHD-LOL4K dataset. Since the inability of most methods to process UHD images directly, the calculation of FLOPs is performed based on the scale of 1024 $\times$ 1024 images (the maximum resolution that these algorithms can process on a single GPU). Our MixNet achieves a better trade-off between model complexity and performance.
Although many learning-based LLIE methods have made significant progress, they always focus on low-resolution images Wei et al. (2018); Zhao et al. (2021); Wu et al. (2022); Xu et al. (2022); Cai et al. (2023). For example, in the LLIE task, the most common dataset used is LOL Wei et al. (2018), which consists of images with resolutions below 1K ($1920\times 1080$). To achieve better results, current state-of-the-art (SOTA) methods Hou et al. (2023); Cai et al. (2023) commonly employ larger models and introduce a significant number of parameters. When faced with UHD images, these SOTA methods often encounter memory overflow and cannot perform full-resolution inference on consumer-grade GPUs. With the prevalence of UHD images, designing a LLIE algorithm tailored explicitly for UHD images becomes increasingly essential.
[…]
In this paper, we propose a simple but effective UHD-LLIE method, called MixNet. Specifically, we design the Feature Mixing Block which mainly contains Global Feature Modulation Layer (GFML), Local Feature Modulation Layer (LFML), and Feed-forward Layer (FFL). GFML is the key component of MixNet, enabling efficient modeling of long-range dependency. The use of self-attention in transformers endows them with powerful long-range dependency modeling capabilities Vaswani et al. (2017). However, this approach comes with significant computational overhead. Inspired by the vanilla MLP-Mixer Tolstikhin et al. (2021), but different from it, GFML encodes the feature maps of images from the perspectives of the width $W$, height $H$, and channel $C$. Through simple dimensional transformation operations, the information encoded in each feature map is associated and fused. It is worth noting that dimension transformation operations do not introduce additional parameters, which means that by permuting the feature maps from different views, GFML achieves efficient modeling of long-range dependency. LFML is designed to capture the local features of images. LFML adaptively calculates a set of parameters for the redistribution of weights in the channel dimension, emphasizing the importance of specific channels. The results of GFML and LFML are fed to FFL which aim to transform features to a compact representation. This way, our MixNet achieves a better trade-off between model complexity and performance. Extensive experiments on both synthetic and real-world datasets demonstrate that MixNet surpasses existing SOTA methods in terms of restoration quality and generalization ability.
Our contributions can be summarized below:
  • We propose a simple but effective UHD-LLIE method called MixNet. MixNet can perform full-resolution inference of UHD images on consumer-grade GPUs. Compared to patch-based strategies, full-resolution inference can avoid the occurrence of boundary artifacts.
  • We develop a novel efficient feature modulation mechanism to capture long-range dependency, which does not need to project the input image from 3D to 2D; it maintains the spatial properties between pixels by modeling directly on 3D space.
  • We evaluate the proposed method quantitatively and qualitatively on synthetic and real-world datasets, and the results show that our MixNet achieves a favorable trade-off between accuracy and model complexity. Furthermore, MixNet shows excellent scalability and flexibility, performing well on other image restoration tasks (e.g., dehazing).
2 Related Work
2.1 Low-Light Image Enhancement
Learning-based methods have been successfully applied to the LLIE task Fu et al. (2022); Guo et al. (2020); Zhou et al. (2022); Wu et al. (2022); Xu et al. (2022). RetinexNet Wei et al. (2018) relies on the Retinex theory and designs a Decom-Net for decomposition and an Enhance-Net for illumination adjustment. Transformers have gained significant traction in computer vision since ViT Dosovitskiy et al. (2020) employs transformers for visual recognition tasks. Restormer Zamir et al. (2022) efficiently achieves image restoration by calculating self-attention at the channel dimension, thereby avoiding significant computational overhead. It demonstrates excellent performance across multiple tasks. Retinexformer Cai et al. (2023) introduces Retinex theory Land (1977) to the transformer and designs an illumination-guided transformer that utilizes illumination representations to direct the modeling of non-local interactions of regions with different lighting conditions. Diffusion models show realistic and detailed image generation through a sequence of denoising refinements. PyDiff Zhou et al. (2023) introduces diffusion models to LLIE to recover realistic details and performs sampling in a pyramid resolution style to get a fast generation. Although the methods above have achieved promising results in the LLIE task, they are not directly applicable to UHD images because the additional pixels of UHD images result in high computational overhead.
Figure 2: The overall architecture of the proposed MixiNet. MixNet first transforms the input low-light image into the feature space using a DownSampler, performs feature extraction using a series of Feature Mixing Blocks (FMBs), and then reconstructs these extracted features using an UpSampler. The FMB mainly contains a Global Feature Modulation Layer (GFML), a Local Feature Modulation Layer (LFML), and a Feed-forward Layer (FFL).
2.2 UHD Image Restoration
UHD image restoration has been an emerging topic in recent years Liang et al. (2021); Zamfir et al. (2023); Guan et al. (2022). In  Zheng et al. (2021a, b), the authors introduced bilateral learning to achieve dehazing of the UHD image and HDR reconstruction of the UHD image. The core idea is to learn the local affine coefficients of the bilateral grid from the low-resolution image and apply them to the full-resolution image. UHD-SFNet Wei et al. (2022) and FourUHD Li et al. (2023) explore underwater UHD image enhancement and UHD low-light image enhancement in the Fourier domain. They are all inspired by the discovery of most luminance information concentrates on amplitudes. Due to the unavoidable need for downsampling in existing UHD image restoration methods, NSEN Yu et al. (2023) proposes a spatial-variant and invertible nonuniform downsampler that adaptively adjusts the sampling rate according to the richness of details. LLFormer Wang et al. (2023) is the first transformer-based method to complete UHD-LLIE task, but like most previous methods, it is unable to perform full-resolution inference on consumer-grade GPUs. Despite the promising results achieved by the aforementioned approaches, there is still room for a favorable trade-off between reconstruction performance and model efficiency.
3 Methodology
MixNet aims to recover ideal UHD normal-light images from low-light images which often exhibit significantly reduced visibility, low contrast, and high levels of noise. We provide the overall pipeline of our method and further details on the critical components of our approach below.
3.1 Overview
An overview of MixNet is shown in Fig. 2. For a given low-light image ${x}\in\mathbb{R}^{H\times W\times 3}$, we first map it to the feature space through a DownSampler to obtain the low-level features ${F}_{0}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent height, width, and channel, respectively. Then, the multiple stacked Feature Mixing Blocks (FMBs) are used to generate finer deep features ${F}_{i}$ from ${F}_{0}$ for normal-light image reconstruction, where a FMB has a Global Feature Modulation Layer (GFML), a Local Feature Modulation Layer (LFML) and a Feed-forward Layer (FFL). Lastly, the sum of the final features ${F}_{f}$ and ${F}_{0}$ is fed to a UpSampler to obtain the restored image ${\hat{y}}$.
Figure 3: The overall architecture of the proposed (a) Global Feature Modulation Layer, (b) Local Feature Modulation Layer, and (c) Feed-forward Layer.
Figure 4: Visualization of dimensional transformation operations. MixNet can capture long-range dependency of features with few parameters by employing simple dimension transformation operations.
3.2 Global Feature Modulation Layer
Recent studies suggest that the notable performance of transformers across diverse tasks stems from their implementation of the key multi-head self-attention (MHSA) mechanism Vaswani et al. (2017). MHSA enables the model with the capability of long-range feature interaction and dynamic spatial weighting, both of which contribute to get promising results. However, this mechanism comes with significant computational overhead. To capture the long-range dependency of features, we ingeniously establish long-range relationships in both spatial and channel dimensions by employing simple dimension transformation operations. This allows for achieving long-range dependency of features by only using a few parameters.
The architecture of the GFML is shown in Fig. 3 (a) and the more details of dimension transformation operations are shown in Fig. 4. Specifically, we first adjust the resolution of the normalized input features, and then perform some dimension transformation operations on them. Given the input feature $F_{in}$, this procedure can be formulated as:
$F_{t}$ $=\operatorname{IP}(F_{in}),$ (1)
$\hat{F}_{t}$ $=\underbrace{\operatorname{Permute}(\operatorname{GELU}(\operatorname{Conv_{1\times 1}}(F_{t})))}_{\times 3},$ (2)
where $F_{t}$ and $\hat{F}_{t}$ are intermediate results. $\operatorname{IP}(\cdot)$ corresponds to the interpolation operation, $\operatorname{Conv_{1\times 1}}(\cdot)$ is a $1\times 1$ convolution, $\operatorname{GELU}(\cdot)$ represents GELU function, $\operatorname{Permute}(\cdot)$ denotes the dimension transformation operation and ×$\times 3$ denotes three operations in sequence. It is worth noting that before executing the last dimension transformation operation, we replace the GELU function with the Sigmoid function.
Afterward, We use interpolation to adjust the feature $\hat{F}_{t}$ to its original resolution to estimate the attention map and adaptively modulate the input $F_{in}$ according to the estimated attention via element-wise product. This process can be written as:
$F_{g}=\operatorname{IP}(\hat{F}_{t})\odot F_{in},$ (3)
where $F_{g}$ are the final output features and $\odot$ represents element-wise product.
3.3 Local Feature Modulation Layer
LFML aims to capture the local feature of images and the architecture of it is shown in Fig. 3 (b). For a set of feature maps, their impact on the final results is different Zhang et al. (2018b), and we hope the model can focus on the important feature maps. We first feed the normalized input features into global average pooling to shift the focus of the model from space dimension to channel dimension, and then feed them into a series of convolutions to get deep features. Finally, these deep features are processed by the sigmoid function to obtain the weight of each channel. The normalized input features adaptively adjust the importance of the channel based on these weights via element-wise product. The LFML can be written as:
$F_{w}$ $=\operatorname{Sigmoid}(\operatorname{Conv}(\operatorname{ReLU}(\operatorname{Conv}(\operatorname{GAP}(F_{in}))))),$ (4)
$F_{l}$ $=F_{w}\odot F_{in},$ (5)
where $F_{in}$ stand for the input features, $F_{w}$ are the channel weights and $F_{l}$ denote the final output features. $\operatorname{Sigmoid}(\cdot)$ represents Sigmoid function, $\operatorname{Conv}(\cdot)$ is a $1\times 1$ convolution and $\operatorname{GAP}(\cdot)$ denotes global average pooling.
3.4 Feed-forward Layer
To transform features into a compact representation, we introduce FFL into our model. As shown in Fig. 3 (c), FFL comprises a $3\times 3$ convolution, a $1\times 1$ convolution, and a GELU function. Within this, the first $3\times 3$ convolution encodes the spatially local contexts and doubles the number of channels of the input features for mixing channels; the later $1\times 1$ convolution reduces the channels back to the original input dimension. In this way, the model learns a more compact representation of features and ignores some unimportant information Tan and Le (2021). The FFL can be expressed as:
$F_{out}=\operatorname{Conv_{1\times 1}(\operatorname{GELU}(\operatorname{Conv}_{3\times 3}((F_{in})))))},$ (6)
where $F_{in}$ and $F_{out}$ are the input features and output features, respectively.
3.5 Feature Mixing Block
Here, we describe our FMB pipeline in general, which can be written as:
$\hat{F}_{in}$ $=\operatorname{Conv}[\operatorname{GFML}(\operatorname{LN}(F_{in}));\operatorname{LFML}(\operatorname{LN}(F_{in}))]+F_{in},$ (7)
$F_{out}$ $=\operatorname{FFL}(\operatorname{LN}(\hat{F}_{in}))+\hat{F}_{in},$ (8)
where $F_{in}$ stand for the input features, $\hat{F}_{in}$ are the intermediate features and $F_{out}$ denote the output features. [;] is concatenation operation, and $\operatorname{Conv}$ respresents the $1\times 1$ convolution.
3.6 Loss Functions
To optimize the weights and biases of the network, we utilize the L1 loss in the RGB color space as the basic reconstruction loss:
$L=\|y-\hat{y}\|_{1},$ (9)
where $y$ and $\hat{y}$ denote the ground truth and restored image, respectively.
Figure 5: Visual quality comparisons. The image can be zoomed in for better visualization. The last row shows the color histogram of the image. Our method has the closest color to GT image.
Table 1: Comparison of quantitative results on UHD-LOL4K and UHD-LL datasets. Best and second best values are indicated with bold text and underlined text respectively.
Type Methods UHD-LOL4K UHD-LL Average Parameter
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
non-UHD Zero-DCE Guo et al. (2020) 17.19 0.850 0.193 17.08 0.663 0.513 17.14 0.757 0.353 79.416K
Zero-DCE++ Li et al. (2021) 15.58 0.835 0.222 16.41 0.630 0.530 16.00 0,733 0.376 10.561K
RUAS Liu et al. (2021) 14.68 0.758 0.274 13.56 0.749 0.460 14.12 0.754 0.367 3.438K
Uformer Wang et al. (2022) 29.99 0.980 0.034 19.28 0.849 0.356 24.64 0.915 0.195 20.628M
Restormer Zamir et al. (2022) 36.91 0.988 0.023 22.25 0.871 0.289 29.58 0.930 0.156 26.112M
UHD LLFormer Wang et al. (2023) 37.33 0.989 0.020 22.79 0.853 0.264 30.06 0.921 0.142 13.152M
UHDFour Li et al. (2023) 36.12 0.990 0.021 26.22 0.900 0.239 31.17 0.945 0.130 17.537M
MixNet 39.22 0.992 0.006 27.54 0.862 0.134 33.38 0.927 0.070 7.770M
4 Experiment
4.1 Implementation Details and Datasets
4.1.1 Implementation Details
We conduct experiments in PyTorch on six NVIDIA GeForce RTX 3090 GPUs. To optimize the network, we employ the Adam optimizer with a learning rate $2\times 10^{-4}$. We randomly crop the full-resolution 4K image to a resolution of $512\times 512$ as the input and perform 300k iterations of training with a batch size of 24. To augment the training data, random horizontal and vertical flips are applied to the input images. The number of FMB and feature channels is set to 8 and 48, respectively. DownSampler and UpSampler are both composed of a sub-pixel convolutional layer Shi et al. (2016) and a $3\times 3$ convolutional layer.
4.1.2 Datasets
UHD-LOL.
UHD-LOL Wang et al. (2023) is a large-scale benchmark dataset containing UHD images in low-light conditions and includes two subsets, UHD-LOL4K and UHD-LOL8K, containing 4K and 8K-resolution images, respectively. In this paper, we use the UHD-LOL4K to verify the effectiveness of MixNet. UHD-LOL4K contains 8,099 image pairs, 5,999 for training and 2, 100 for testing.
UHD-LL.
UHD-LL Li et al. (2023) is a real-world low-noise/normal-clear paired image dataset that contains 2,150 pairs of 4K UHD data saved in 8bit sRGB format. It has been split into two parts: 2,000 pairs for training and 115 pairs for testing.
LOL.
In addition to the aforementioned UHD datasets, we also evaluate our method on the LOL dataset Wei et al. (2018) which is the commonly used dataset in the LLIE task. LOL has 500 image pairs and the training and testing sets are split in proportion to 485:15.
Figure 6: Visual quality comparisons. The image can be zoomed in for improved visualization. Our method can effectively remove the noise and reconstruct well-exposed image details.
4.2 Comparisons with the State-of-the-art Methods
4.2.1 Quantitative Results
We employ three well-known image quality assessment metrics, PSNR, SSIM Wang et al. (2004) and LPIPS (Alex version) Zhang et al. (2018a) to quantify the performance of different methods. PSNR and SSIM are calculated along the RGB channels. For UHD datasets, we compare our method with Zero-DCE Guo et al. (2020), Zero-DCE++ Li et al. (2021), RUAS Liu et al. (2021), Uformer Wang et al. (2022), Restormer Zamir et al. (2022), LLFormer Wang et al. (2023) and UHDFour Li et al. (2023). It is worth noting that most methods cannot directly perform full-resolution inference on UHD images. Therefore, for these methods, we adopt a strategy of dividing the images into multiple patches without overlapping for interference, and then the results are stitched together. For the LOL dataset, in addition to the mentioned methods, we also compare our MixNet with methods RetinexNet Wei et al. (2018), KinD Zhang et al. (2019), DSLR Lim and Kim (2020), RUAS Liu et al. (2021), ELGAN Jiang et al. (2021) and Retinexformer Cai et al. (2023).
In Tab. 1, we report the quantitative comparison results among different methods on the UHD dataset. When compared with the previous SOTA method FourUHD, our MixNet gains by an average of 2.21dB on the UHD-LOL4K and UHD-LL datasets. Compared to other methods specifically designed for UHD images, our MixNet achieves impressive results with few parameters. Our method also performs well on the LOL dataset and the results are reported in Tab. 2. It is worth noting that the resolution of LOL images is relatively low. Therefore, we replace the sub-pixel convolutional layer in DownSampler and UpSampler with the Residual Block of comparable parameters.
All these results clearly suggest the outstanding effectiveness and efficiency advantage of our MixNet.
4.2.2 Qualitative Results
In addition to the quantitative evaluations, we also provide qualitative comparisons of the proposed method. Figs. 5 and  6 present the visual comparisons. MixNet can effectively enhance the poor visibility and low contrast or low-light regions, reliably remove the noise without introducing spots and artifacts, and robustly preserve the color.
Table 2: Comparison of quantitative results on LOL dataset. Best and second best values are indicated with bold text and underlined text respectively.
Type Methods PSNR SSIM
non-UHD RetinexNet Wei et al. (2018) 16.77 0.425
KinD Zhang et al. (2019) 17.65 0.772
DSLR Lim and Kim (2020) 14.98 0.596
Z_DCE Guo et al. (2020) 14.86 0.562
Z_DCE++ Li et al. (2021) 14.75 0.518
RUAS Liu et al. (2021) 16.40 0.503
ELGAN Jiang et al. (2021) 17.48 0.652
Uformer Wang et al. (2022) 18.55 0.721
Restormer Zamir et al. (2022) 22.37 0.816
Retinexformer Cai et al. (2023) 25.16 0.845
UHD UHDFour Li et al. (2023) 23.09 0.871
LLFormer Wang et al. (2023) 23.65 0.816
MixNet 24.21 0.860
4.3 Ablation Study
We further conduct extensive ablation studies to better understand and evaluate each component in the proposed MixNet. For fair comparisons with the designed baselines, we implement all experiments with the same setting on UHD-LOL4K dataset. We gradually replace the proposed module with the Residual Block (ResBlock) He et al. (2016) of comparable parameters, and the experimental results in Tab. 4.
Based on all the results in the Tab. 4, all the key designs contribute to the best performance of the full model. Without the GFML (#a, #c, #d, #f), the model shows the most severe performance degradation. This aligns with our hypothesis, as the excellent long-range dependency modeling capability brought by GMFL is one of the key factors contributing to the outstanding performance of MixNet. And the result suggests the importance of long-range dependency modeling among features. From the results of #g and #e, we find that the impact of FFL on performance is higher than that of GFML. We believe that this might be the ability of the FFL to learn compact feature representations, which the ResBlock lacks. And any CNN is capable of learning local features. From the results of #c and #d, we observe that when using the block we proposed alone, LFML has a higher impact on the results than FFL. This may be because when the feature representation is not effective enough (not from GFML and LFML), learning its compact representation has little effect. We also provide visualizations of the feature maps corresponding to #e, #f, #g and #h. As shown in Fig. 7, it is evident that our method reconstructs feature maps with sharper textures.
Table 3: Comparison of quantitative results on O-Haze dataset. Best and second best values are indicated with bold text and underlined text respectively.
Task Methods PSNR SSIM
non-UHD DCP He et al. (2010) 14.54 0.333
BCCR Meng et al. (2013) 15.20 0.361
CAP Zhu et al. (2015) 15.35 0.342
NLD Berman et al. (2016) 14.76 0.396
DehazeNet Cai et al. (2016) 12.39 0.359
AodNet Li et al. (2017) 15.87 0.581
GridNet Liu et al. (2019) 17.44 0.603
PMSNet Chen et al. (2019) 16.99 0.597
MSBDN Dong et al. (2020) 17.46 0.635
UHD MGBL Zheng et al. (2021a) 18.72 0.614
NSEN Yu et al. (2023) 19.23 0.672
MixNet 21.46 0.650
4.4 Extension Experiments
Although we primarily focus on the UHD-LLIE task in this paper, it is important to note that we do not introduce specific prior knowledge related to low-light conditions. Therefore, our method can be considered as a general approach that performs well not only in UHD-LLIE task but also in various UHD image restoration task, such dehazing.
4.4.1 Datasets
We conduct dehazing experiments on the O-HAZE dataset which has 45 paired hazy and haze-free images. Following the prior work Yu et al. (2023), we randomly choose 35 images for training and the others for testing.
4.4.2 Results
We compare the proposed MixNet with several image dehazing approaches Berman et al. (2016); He et al. (2010); Meng et al. (2013); Zhu et al. (2015); Li et al. (2017); Dong et al. (2020); Liu et al. (2019); Cai et al. (2016); Chen et al. (2019); Zheng et al. (2021a); Yu et al. (2023). Results are reported in Tab. 3. MixNet achieves a substantial gain of over 2 dB compared to NSEN in PSNR on the UHD image dehazing task.
Table 4: Ablation study of proposed blocks. In order to ensure consistency in parameter magnitude, we gradually replace the proposed module with Residual Block of comparable parameters.
# GFML LFML FFL PSNR SSIM Parameter
(a) 30.73 0.968 7.178M
(b) 34.16 0.985 7.475M
(c) 33.93 0.984 6.882M
(d) 32.52 0.976 7.769M
(e) 36.44 0.990 8.066M
(f) 34.47 0.987 7.474M
(g) 35.80 0.988 7.179M
(h) 39.22 0.992 7.770M
Figure 7: Visualizations of the feature maps before the UpSampler. The image used here is the low-light image shown in Fig. 6. Our method reconstructs feature maps with sharper textures.
5 Conclusion
In this paper, we proposed a simple yet efficient framework called MixNet to solve the efficient UHD-LLIE problem. MixNet has a series of FMBs which mainly contain a GFML, LFML, and FFL. GFML explored long-range dependency modeling upon the simple dimensional transformation operation. By permuting the feature maps, GFML associates features from different views. LFML and FFL are designed to capture local features and transform features into a compact representation, respectively. Extensive experiments on both synthetic and real-world datasets showed that the proposed MixNet is more efficient than state-of-the-art methods while achieving competitive performance.